once kg of air expand at a constant temperature from a pressure of 820 kpa and a volume of 3 m³ to a pressure of 220 kpa. determine. a. change of Internal Energy b. change of Enthalpy

1
Expert's answer
2022-02-18T08:22:44-0500

a)

"\Delta T=0,\implies U=0~ J,"

b)

"H=U+p(V_2-V_1)=p(V_2-V_1)=1.8~MJ."